OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch radius A = sq. in.

OpenStudy (anonymous):

what radius, circum or in?

OpenStudy (ledah):

For equilateral triangles, when s = the length of a side, the area equals s squared times the square root of 3 over 4 (area = s^2(sqrt(3)/4)). So when s = 6, the area = 36*sqrt(3)/4 = 9*sqrt(3). Hope this helps better your understanding.

OpenStudy (anonymous):

But here s is not 6.

OpenStudy (anonymous):

Radius of a triangle?? How?

OpenStudy (ledah):

ah, true. but i only did an example...

OpenStudy (anonymous):